OpenStudy (anonymous):

Solve the equation? t(3t+1)=0

OpenStudy (anonymous):

t=0 3t+1=0 3t=-1 t=-1/3 t=0 , t=-1/3

OpenStudy (anonymous):

so the set is (0,-1/3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you